You were hired as a consultant to Quigley Company, whose target capital structure is 35% debt, 10% preferred, and 55% common equity. The interest rate on new debt is 6.50%, the yield on the preferred is 6.00%, the cost of retained earnings is 14.75%, and the tax rate is 40%. The firm will not be issuing any new stock. What is Quigley's WACC? Round final answer to two decimal places. Do not round your intermediate calculations.

a.

8.36%

b.

9.17%

c.

12.19%

d.

8.87%

e.

10.08%
Answer #1

After-tax cost of debt=6.5*(1-tax rate)

=6.5*(1-0.4)=3.9%

WACC=Respective costs*Respective weight

=(3.9*0.35)+(0.10*6)+(0.55*14.75)

=10.08%(Approx).